OpenStudy (anonymous):

Simplify, showing your steps. Tell the property you used in each step. 1/4(1*4)+(-4+4) Simplify, showing your steps. Tell the property you used in each step. 1/4(1*4)+(-4+4) @Mathematics

OpenStudy (anonymous):

\[\frac{1}{4}(1 * 4)+ (-4 + 4)\] \[= \frac{1}{4}(1*4) + 0\ \ \ \ \text{Additive Inverse property}\] \[= \frac{1}{4}(1*4) \ \ \ \ \text{Additive Identity property}\] \[= \frac{1}{4}(4)\ \ \ \ \text{Multiplicative Identity property}\] \[= 1\ \ \ \ \text{Multiplicative Inverse property}\]

OpenStudy (anonymous):

Thank you (: do you mind helping me with my others?